Please help me solve for x and y in the following linear system.
x-y=5
2x-y=13

Thank you.

Answer by mananth(16946)

x-y=5.................1
2x-y=13...............2
..
multiply 1 by -1
-x+y=-5
add it to 2
x=8
plug value of x in (1)
we get y= 3
...
m.ananth@hotmail.ca